# Visually Robust Adversarial Imitation Learning from Videos with Contrastive Learning

### v1

@@ 5 Experiments
The results of these experiments are summarized in Table 1, which includes results for C-LAIfO and its various configurations. All the learning curves are provided in Appendix D. In Table 1, C-LAIfO is implemented as described in Section 4, with the data augmentation function $\text{aug}(\cdot)$ defined as a brightness transformation. In BYOL-LAIfO, we retain the design identical to C-LAIfO except for replacing the InfoNCE loss in (5) with BYOL [4]. In C-LAIfO w/o $Q$ backprop, we disable gradient backpropagation from $Q_{\bm{\psi}k}$ to $\phi_{\bm{\delta}}$ in (3). Lastly, in C-LAIfO full aug and C-LAIfO w/o aug, we respectively modify the data augmentation function $\text{aug}(\cdot)$ to include full augmentation (brightness, color, and geometric transformations, as detailed in Appendix 5) and no augmentation. Notably, without backpropagating the gradient from $Q_{\bm{\psi}_{k}}$ to $\phi_{\bm{\delta}}$, C-LAIfO fails to solve the imitation task even in the simplest mismatch scenario in (2(b)). This result demonstrates the importance of this step for embedding goal-completion information in $\mathcal{Z}$. Conversely, without a well-designed $\text{aug}(\cdot)$, C-LAIfO can solve the imitation problem with minor visual mismatch but struggles with more challenging visual distractors. These results highlight the critical role of properly defining $\text{aug}(\cdot)$ for efficient visual generalization in $\mathcal{Z}$. General $\text{aug}(\cdot)$ might still achieve this goal, but much less efficiently. Finally, the superior performance of the InfoNCE loss in (5) compared to BYOL is particularly evident in handling the most difficult mismatches in (2(e)) and (2(f)).
Figure 2: Different environments used for the ablation study in Table 1. The environment in (2(a)) represents the source-POMDP used to collect expert data, while (2(b))–(2(f)) are different target-POMDPs. In these experiments, visual mismatch is introduced by varying the light intensity, with the degree of mismatch increasing linearly from (2(b)) to (2(f)).
Table 1: Summary of the ablation experiments. Refer to Fig. 2 for a visual representation of the different environments. We use DDPG [44] to train experts in a fully observable setting and collect $100$ episodes of expert data. All algorithms are trained for $10^{6}$ steps, and the learned policies are evaluated based on average return over $10$ episodes. We report the mean and standard deviation of the final return over $6$ seeds and highlight the best performance. The experiments C-LAIfO w/o $Q$ backprop, C-LAIfO full aug, and C-LAIfO w/o aug are only conducted on the settings in (2(b)) and (2(c)) due to their low performance on these two easier settings.
Source Env (2(a)), Performance = $950$
Target Env (2(b)) (2(c)) (2(d)) (2(e)) (2(f))
C-LAIfO $936\pm 33.2$ $808\pm 269$ $526\pm 334$ $768\pm 231$ $895\pm 36.6$
BYOL-LAIfO $906\pm 76.1$ $707\pm 337$ $696\pm 333$ $142\pm 124$ $168\pm 139$
C-LAIfO w/o $Q$ backprop $114\pm 25.0$ $48.7\pm 8.7$
C-LAIfO full aug $816\pm 209$ $96.8\pm 53.4$
C-LAIfO w/o aug $641\pm 327$ $113\pm 25.3$
Visual Imitation from Observations with mismatch
In this section, we test C-LAIfO in the V-IfO setting with different types of mismatches (cf. Fig. 3) and compare it with three baselines: LAIfO [1] and PatchAIL [2], both equipped with the same $\text{aug}(\cdot)$ used for C-LAIfO, and DisentanGAIL [6]. The results, summarized in Table 2, demonstrate that C-LAIfO successfully addresses the V-IfO with mismatch problem, achieving superior performance compared to all the baselines across the proposed mismatches. All the learning curves are provided in Appendix D. For the Light experiment, $\text{aug}(\cdot)$ is defined as a brightness transformation; while for the others it is defined as a color transformation. Details on $\text{aug}(\cdot)$ are provided in Appendix 5. Furthermore, to assess whether C-LAIfO achieves domain invariance in the feature space $\mathcal{Z}$, we perform PCA and t-SNE analyses on the latent space $\mathcal{Z}$ learned by different algorithms during training. Specifically, focusing on the Light experiment in Table 2, Fig. 4(a) compares the latent space $\mathcal{Z}$ learned by LAIfO in [1], C-LAIfO, and LAIfO with data augmentation. Additionally, Fig. 4(b) investigates whether the latent space learned by C-LAIfO during the Full experiment can generalize to the unseen environment depicted in Fig. 3(g). The results of this analysis are shown in Fig. 4, which includes the PCA visualization. For the t-SNE visualization and additional details refer to Appendix E.
Figure 3: Different environments used for the experiments in Table 2 and the PCA in Fig. 4.
Table 2: Summary of the experiments for the mismatches in Fig. 3. The Light experiment consists in (3(b)) as source-POMDP and (3(a)) as target-POMDP. The Body, Floor, Background, and Full experiments have (3(b)) as target-POMDP and respectively (3(c)), (3(d)), (3(e)), and (3(f)) as source-POMDP. We use DDPG [44] to train experts in a fully observable setting and collect $100$ episodes of expert data. We train all the algorithms for $10^{6}$ steps, except for the Full mismatch, where we train for $2\times 10^{6}$ steps. We evaluate the learned policies based on the average return over $10$ episodes. We report mean and standard deviation over $10$ seeds and highlight the best performance.
Light Body Floor Background Full
C-LAIfO $895\pm 36.6$ $779\pm 199$ $486\pm 259$ $702\pm 164$ $509\pm 235$
LAIfO [1] w/ data aug $64.6\pm 62.9$ $714\pm 222$ $433\pm 304$ $315\pm 329$ $206\pm 210$
DisentanGAIL [6] $28.1\pm 8.7$ $337\pm 215$ $75.6\pm 59.8$ $66.8\pm 53.9$ $30.6\pm 13.2$
PatchAIL [2] w/ data aug $18.4\pm 5.6$ $597\pm 247$ $14.1\pm 2.4$ $29.9\pm 10.3$ $122\pm 66.6$
(a) PCA results for the Light experiment in Table 2.
@@
(b) PCA results on C-LAIfO for the Full experiment in Table 2 and the unseen environment in Fig. 3(g).
Figure 4: In Fig. 4(a) we compare the latent space $\mathcal{Z}$ learned by LAIfO, C-LAIfO, and LAIfO with data augmentation in the Light setting from Table 2. Specifically, we define source and target as the observations generated by an optimal policy in the source and target POMDPs, respectively. Similarly, source random and target random are observations generated by random policies. These observations are processed with the encoder $\phi_{\bm{\delta}}:\mathcal{X}^{d}\to\mathcal{Z}$ trained using the respective algorithms. We perform PCA on this set of latent variables $z_{1:T}$ and plot the first two principal components. The results show that C-LAIfO is the only algorithm capable of filtering out visual distractors and clustering together data points with the same goal-completion information. In Fig. 4(b), we focus on C-LAIfO and test for generalization to the unseen environment in Fig. 3(g). We train $\phi_{\bm{\delta}}$ on the Full setting from Table 2 and perform PCA as described. The results in the unseen setting match those in the Full experiment, indicating that $\phi_{\bm{\delta}}$ trained on (3(f)) can successfully generalize to (3(g)).
Improving RL using expert videos
In the following section, we examine a scenario where the reward $r_{\bm{\chi}}$ learned by C-LAIfO can be combined with a sparse reward $\mathcal{R}$ that the agent collects through interaction with the environment. Consequently, the RL problem aims to maximize the total reward $\mathcal{R}_{\text{tot}}=\mathcal{R}(s_{t},a_{t})+r_{\bm{\chi}}(z_{t},z_{t+1})$ with $r_{\bm{\chi}}$ learned through the AIL step in (2). We conduct experiments on the Adroit platform for dynamic dexterous manipulation [7] and compare C-LAIfO with DrQ-v2 [45], an RL-only baseline. In these experiments, both C-LAIfO and DrQ-v2 use an encoder to process pixel observations and extract embeddings in $\mathcal{Z}$, which are then concatenated with the robot sensory observations. Importantly, expert’s sensory observations are not used in the imitation problem as we only assume access to videos of experts. We denote our approach as RL+C-LAIfO since it maximizes $\mathcal{R}_{\text{tot}}$ rather than only $r_{\bm{\chi}}$ alone as in the standard imitation learning problem. The results, summarized in Table 3, show that DrQ-v2 fails to learn in two out of three environments. In contrast, C-LAIfO successfully leverages expert videos with visual mismatch to facilitate learning. Note that the goal of these experiments is not merely to compare RL+C-LAIfO with DrQ-v2 but to demonstrate how C-LAIfO can leverage expert videos, even with visual mismatch, to enable learning in challenging robotic tasks. We provide all the learning curves in Appendix D.
Figure 5: Adroit environments used for the experiments in Table 3.
@@
Table 3: Summary of the experiments in Fig. 5. The Door-Light and Door-Color experiments consider (5(a)) as the source-POMDP and (5(b)) and (5(c)) as the respective target-POMDPs. Similarly, the Hammer-Light and Hammer-Color experiments consider (5(d)) as source-POMDP and (5(e)) and (5(f)) as the respective target-POMDPs, while the Pen-Light and Pen-Color experiments consider (5(g)) as source-POMDP and (5(h)) and (5(i)) as the respective target-POMDPs. We use the VRL3 pipeline in [46] to train expert policies and collect $100$ episodes of expert data. All algorithms are trained for $4\times 10^{6}$ steps. We evaluate the learned policies based on the average return over $10$ episodes. We report the mean and standard deviation over $10$ seeds and highlight the best performance.
@@
DrQ-v2 [45] $-2.4\pm 0.0$ $-1.8\pm 2.2$ $53\pm 9.0$
RL+C-LAIfO $106\pm 75$ $96\pm 80$ $181\pm 4.0$ $103\pm 86$ $59\pm 12$ $59\pm 4.8$
6 Conclusion
@@
In this work, we analyze the V-IfO problem in the presence of visual mismatches and propose a novel algorithm named C-LAIfO as an effective solution. Through comprehensive ablation studies, we provide insights into our design choices and demonstrate the superior performance of our approach compared to a range of baselines in imitation from videos under various mismatch scenarios. Furthermore, we illustrate how C-LAIfO effectively utilizes expert videos with visual mismatches to facilitate learning in challenging hand manipulation tasks characterized by sparse rewards and realistic visual inputs.
Limitations and future work
@@
A main limitation of the current approach is given by the reliance of C-LAIfO on a well-designed, possibly mismatch-informed, data augmentation function. As illustrated in our ablations in Section 5, general augmentation can lead to poor performance or can remarkably reduce the algorithmic sample efficiency. Furthermore, it can be challenging to design effective augmentations for certain types of mismatches. To address this problem, exploring generative models for automatic data augmentation represents an interesting research direction. Generative models could produce diverse, mismatch-informed augmentations, potentially overcoming the limitations of manually designed strategies. Alternatively, investigating different auxiliary losses that are less reliant on augmentation techniques represents another interesting direction. Finally, future work will be devoted to go beyond simulated environments and test our algorithms on hardware in real-world scenarios.
7 Acknowledgements
Vittorio Giammarino and Ioannis Ch. Paschalidis were partially supported by the NSF under grants IIS-1914792, CCF-2200052, and ECCS-2317079, by the ONR under grant N00014-19-1-2571, by the DOE under grant DE-AC02-05CH11231, by the NIH under grant UL54 TR004130, and by Boston University. James Queeney was exclusively supported by Mitsubishi Electric Research Laboratories.
Appendix A Analysis
@@
Appendix B Pseudo-code and hyperparameters
@@
Table 4: Hyperparameter values for C-LAIfO experiments.
@@
Appendix C Data augmentation
The operations used in our data augmentation functions for the experiments in Section 5 are summarized in Table 5. Fig. 6 shows an example of color augmentation as implemented for the experiments in Table 2, where we randomly perform all the operations in the color transformations row in Table 5. On the other hand, Fig. 7 shows only a brightness transformation as implemented in the ablation study in Table. 1. Full augmentation in Table 1 performs all the operations in Table 5. For additional details refer to our code [^1] https://github.com/VittorioGiammarino/C-LAIfO.
Table 5: Operations used in our data augmentation function.
@@ Appendix D Learning curves
Fig. 8, Fig. 9, and Fig. 10 show the learning curves for the results in Table 1, Table 2, and Table 3, respectively.
@@
Figure 8: Learning curves for the results in Table 1. Plots show the average return per episode as a function of training steps. The environment in (8(a)) represents the source-POMDP used to collect expert data, while (8(b))–(8(f)) are different target-POMDPs. In these experiments, visual mismatch is introduced by varying the light intensity, with the degree of mismatch increasing linearly from (8(b)) to (8(f)).
@@
Figure 9: Learning curves for the results in Table 2. Plots show the average return per episode as a function of training steps. The Light experiment consists in (9(b)) as source-POMDP and (9(a)) as target-POMDP. The Body, Floor, Background, and Full experiments have (9(b)) as target-POMDP and respectively (9(c)), (9(d)), (9(e)), and (9(f)) as source-POMDP.
@@
Figure 10: Learning curves for the results in Table 3. Plots show the average return per episode as a function of training steps. The Door-Light and Door-Color experiments consider (10(a)) as the source-POMDP and (10(b)) and (10(c)) as the respective target-POMDPs. Similarly, the Hammer-Light and Hammer-Color experiments consider (10(d)) as source-POMDP and (10(e)) and (10(f)) as the respective target-POMDPs, while the Pen-Light and Pen-Color experiments consider (10(g)) as source-POMDP and (10(h)) and (10(i)) as the respective target-POMDPs.
Appendix E PCA and t-SNE analysis
@@
In Fig. 11 we compare the latent space $\mathcal{Z}$ learned by LAIfO, C-LAIfO, and LAIfO with data augmentation in the Light setting from Table 2 and Fig. 8. Specifically, we define source and target as the observations generated by an optimal policy in the source and target POMDPs, respectively. Similarly, source random and target random are observations generated by random policies. These observations are processed with the encoder $\phi_{\bm{\delta}}:\mathcal{X}^{d}\to\mathcal{Z}$ trained using the respective algorithms. We perform PCA and t-SNE on this set of latent variables $z_{1:T}$ and plot the first two principal components. The results show that C-LAIfO is the only algorithm capable of filtering out visual distractors and clustering together data points with the same goal-completion information. In Fig. 12, we focus on C-LAIfO and test for generalization to the unseen environment in Fig. 12(c). We train $\phi_{\bm{\delta}}$ on the Full setting from Table 2 and Fig. 9 and perform PCA and t-SNE as described. In Fig. 12, the Full figures have (12(a))-(12(b)) as source-POMDP and (12(e))-(12(f)) as target-POMDP. Similarly the Full-unseen figures have (12(c))-(12(d)) as source-POMDP and (12(e))-(12(f)) as target-POMDP. The results in the Full-unseen setting match those in the Full experiment, indicating that $\phi_{\bm{\delta}}$ trained on (12(a)) can successfully generalize to (12(c)).
@@
Figure 11: PCA and t-SNE visualizations for the Light experiment in Table 2.
@@
Figure 12: PCA and t-SNE for the Full experiment in Table 2. Fig. 12(a) and Fig. 12(b) denote the source environment for the Full experiment (left figures in Fig. 12). Fig. 12(c) and Fig. 12(d) denote the source environment for the Full-unseen experiment (right figures in Fig. 12).

### v2

@@ V-A Ablation study
These results are summarized in Table I, which includes results for C-LAIfO and its various configurations. All the learning curves are provided in the Arxiv version of the paper. In Table I, C-LAIfO is implemented as described in Section IV, with the data augmentation function $\text{aug}(\cdot)$ defined as a brightness transformation. In BYOL-LAIfO, we retain the design identical to C-LAIfO except for replacing the InfoNCE loss in (5) with BYOL [4]. In C-LAIfO w/o $Q$ backprop, we disable gradient backpropagation from $Q_{\bm{\psi}k}$ to $\phi_{\bm{\delta}}$ in (3). Lastly, in C-LAIfO full aug and C-LAIfO w/o aug, we respectively modify the data augmentation function $\text{aug}(\cdot)$ to include full augmentation (brightness, color, and geometric transformations, as detailed in the Arxiv version of the paper) and no augmentation. Notably, without backpropagating the gradient from $Q_{\bm{\psi}_{k}}$ to $\phi_{\bm{\delta}}$, C-LAIfO fails to solve the imitation task even in the simplest mismatch scenario. This result demonstrates the importance of this step for embedding goal-completion information in $\mathcal{Z}$. Similar considerations can be done for the design of $\text{aug}(\cdot)$ where C-LAIfO struggles to efficiently solve the task both when the augmentation is too generic and when it is absent. These results highlight the critical role of properly defining $\text{aug}(\cdot)$ for efficient visual generalization in $\mathcal{Z}$. Finally, the superior performance of the InfoNCE loss in (5) compared to BYOL is evident in handling the hardest mismatch in Table I.
V-B Visual Imitation from Observations with mismatch
Figure 3: Different environments used for the experiments in Table II and the PCA in Fig. 4 and 5.
TABLE II: Summary of the experiments for the mismatches in Fig. 3. The Light experiment consists in (3(b)) as source-POMDP and (3(a)) as target-POMDP. The Full experiments have (3(b)) as target-POMDP and (3(c)) as source-POMDP. We train all the algorithms in the Light mismatch for $10^{6}$ steps and in the Full mismatch for $2\times 10^{6}$ steps.
Light Full
C-LAIfO $895\pm 36.6$ $509\pm 235$
LAIfO [1] w/ data aug $64.6\pm 62.9$ $206\pm 210$
DisentanGAIL [5] $28.1\pm 8.7$ $30.6\pm 13.2$
PatchAIL [2] w/ data aug $18.4\pm 5.6$ $122\pm 66.6$
Figure 4: PCA results for the Light experiment in Table II.
@@ V-B Visual Imitation from Observations with mismatch
Figure 5: PCA results on C-LAIfO for the Full experiment in Table II and the unseen environment in Fig. 3(d).
Figure 6: Adroit environments used for the experiments in Table III.
@@
TABLE III: Summary of the experiments in Fig. 6. The Door-Light and Door-Color experiments consider (6(a)) as the source-POMDP and (6(b)) and (6(c)) as the respective target-POMDPs. Similarly, the Hammer-Light and Hammer-Color experiments consider (6(d)) as source-POMDP and (6(e)) and (6(f)) as the respective target-POMDPs, while the Pen-Light and Pen-Color experiments consider (6(g)) as source-POMDP and (6(h)) and (6(i)) as the respective target-POMDPs. We use the VRL3 pipeline in [40] to train expert policies and collect $100$ episodes of expert data. For these experiments, all the algorithms are trained for $4\times 10^{6}$ steps and we report mean and standard deviation over 10 random seeds.
@@
RL+LAIfO [1] $106\pm 75$ $96\pm 80$ $181\pm 4.0$ $103\pm 86$ $59\pm 12$ $59\pm 4.8$
RL+C-LAIfO $165\pm 5.8$ $160\pm 12$ $183\pm 1.4$ $178\pm 11$ $64\pm 5.6$ $62\pm 6.9$
In this section, we test C-LAIfO in the V-IfO setting with different types of mismatches (cf. Fig. 3) and compare it with three baselines: LAIfO [1] and PatchAIL [2], both equipped with the same $\text{aug}(\cdot)$ used for C-LAIfO, and DisentanGAIL [5]. The results, summarized in Table II, demonstrate that C-LAIfO successfully addresses the V-IfO with mismatch problem, achieving superior performance compared to all the baselines across the proposed mismatches. All the learning curves are provided in the Arxiv version of the paper. For the Light experiment, $\text{aug}(\cdot)$ is defined as a brightness transformation; while for the others it is defined as a color transformation. Details on $\text{aug}(\cdot)$ are provided in the Arxiv version of the paper. Furthermore, to assess whether C-LAIfO achieves domain invariance in the feature space $\mathcal{Z}$, we perform PCA on the latent space $\mathcal{Z}$ learned by different algorithms during training. In Fig. 4 we compare the latent space $\mathcal{Z}$ learned by LAIfO, C-LAIfO, and LAIfO with data augmentation in the Light setting from Table II. Specifically, we define source and target as the observations generated by an optimal policy in the source and target POMDPs, respectively. Similarly, source random and target random are observations generated by random policies. These observations are processed with the encoder $\phi_{\bm{\delta}}:\mathcal{X}^{d}\to\mathcal{Z}$ trained using the respective algorithms. We perform PCA on this set of latent variables $z_{1:T}$ and plot the first two principal components. The results show that C-LAIfO is the only algorithm capable of filtering out visual distractors and clustering together data points with the same goal-completion information. In Fig. 5, we focus on C-LAIfO and test for generalization to the unseen environment in Fig. 3(d). We train $\phi_{\bm{\delta}}$ on the Full setting from Table II and perform PCA as described. The results in the unseen setting match those in the Full experiment, indicating that $\phi_{\bm{\delta}}$ trained on (3(c)) can successfully generalize to (3(d)). Refer to the Arxiv version of the paper also for the t-SNE visualization and additional details.
V-C C-LAIfO for dexterous manipulation
In the following section, we evaluate our algorithm on a series of challenging robotic manipulation tasks from the Adroit platform for dynamic dexterous manipulation [6]. These experiments demonstrate how the reward $r_{\bm{\chi}}$, learned by C-LAIfO from expert videos, can be effectively combined with a sparse reward $\mathcal{R}$, collected by the agent through interaction with the environment, to enhance learning efficiency. The RL problem, therefore, aims to maximize the total reward $\mathcal{R}_{\text{tot}}=\mathcal{R}(s_{t},a_{t})+r_{\bm{\chi}}(z_{t},z_{t+1})$, where $r_{\bm{\chi}}$ is learned through the AIL step in (2). This approach is particularly relevant for robotic tasks, where sparse rewards are often the most feasible option in real-world settings. However, relying solely on sparse rewards can make learning highly challenging and inefficient [41]. In this context, leveraging expert videos can significantly enhance efficiency. We compare C-LAIfO with the standard LAIfO algorithm [1], which does not explicitly address the visual mismatch between source and target POMDPs. Both C-LAIfO and LAIfO utilize an encoder to process pixel observations and extract embeddings in $\mathcal{Z}$, which are then concatenated with robot sensory observations. Notably, the expert’s sensory observations are not used in the imitation process, as we only assume access to expert videos. Our approach, denoted as RL+C-LAIfO (or RL+LAIfO), seeks to maximize $\mathcal{R}_{\text{tot}}$, rather than just $r_{\bm{\chi}}$, as in the standard imitation learning problem. The results, summarized in Table III, show that C-LAIfO more effectively leverages expert videos with visual mismatches to facilitate learning when compared to LAIfO. This demonstrates the potential of our approach to enable learning in challenging robotic tasks by utilizing a minimal form of supervision, relying solely on expert videos. All learning curves are provided in the Arxiv version of the paper.
VI Conclusion
@@
In this work, we analyze the V-IfO problem with visual mismatches and propose a novel algorithm named C-LAIfO as an effective solution. Through comprehensive ablation studies, we provide insights into our design and demonstrate the superior performance of our approach compared to a range of baselines in imitation from videos under various mismatch scenarios. Furthermore, we illustrate how C-LAIfO effectively utilizes expert videos with visual mismatches to facilitate learning in challenging hand manipulation tasks characterized by sparse rewards and realistic visual inputs.
@@
A main limitation of the current approach is given by the reliance of C-LAIfO on a well-designed, possibly mismatch-informed, data augmentation function. As illustrated in our ablations in Section V, general augmentation can lead to poor performance or can remarkably reduce the algorithmic sample efficiency. Furthermore, it can be challenging to design effective augmentations for certain types of mismatches. To address this problem, exploring generative models for automatic data augmentation represents an interesting research direction. Generative models could produce diverse, mismatch-informed augmentations, potentially overcoming the limitations of manually designed strategies. Alternatively, investigating different auxiliary losses that are less reliant on augmentation techniques represents another interesting direction. Finally, future work will be devoted to go beyond simulated environments and test our algorithms on hardware in real-world scenarios.
Supplementary material
VI-A Analysis
@@
VI-B Pseudo-code and hyperparameters
@@
TABLE IV: Hyperparameter values for C-LAIfO experiments.
@@
VI-C Data augmentation
The operations used in our data augmentation functions for the experiments in Section V are summarized in Table V. Fig. 7 shows an example of color augmentation as implemented for the experiments in Table II, where we randomly perform all the operations in the color transformations row in Table V. On the other hand, Fig. 8 shows only a brightness transformation as implemented in the ablation study in Table. I. Full augmentation in Table I performs all the operations in Table V. For additional details refer to our code [^1] https://github.com/VittorioGiammarino/C-LAIfO.
TABLE V: Operations used in our data augmentation function.
@@ VI-D Learning curves
Fig. 9, Fig. 11, and Fig. 12 show the learning curves for the results in Table I, Table II, and Table III, respectively. Fig. 10 shows the effect of randomized data augmentation when used in the AIL step in (2).
@@
Figure 9: Learning curves for the results in Table I. Plots show the average return per episode as a function of training steps. The environment in (9(a)) represents the source-POMDP used to collect expert data, while (9(b))–(9(c)) are different target-POMDPs. In these experiments, visual mismatch is introduced by varying the light intensity in the target-POMDPs.
Figure 10: Experiments on the use of data augmentation in AIL in (2). Plots show the average return per episode as a function of training steps. The environment in (9(a)) represents the source-POMDP used to collect expert data, while (9(c)) the target-POMDPs. These experiments show how randomized data augmentation used during AIL triggers instability.
@@
Figure 11: Learning curves for the results in Table II. Plots show the average return per episode as a function of training steps. The Light experiment consists in (11(b)) as source-POMDP and (11(a)) as target-POMDP. The Full experiment considers (11(b)) as target-POMDP and (11(c)) as source-POMDP.
@@
Figure 12: Learning curves for the results in Table III. Plots show the average return per episode as a function of training steps. The Door-Light and Door-Color experiments consider (12(a)) as the source-POMDP and (12(b)) and (12(c)) as the respective target-POMDPs. Similarly, the Hammer-Light and Hammer-Color experiments consider (12(d)) as source-POMDP and (12(e)) and (12(f)) as the respective target-POMDPs, while the Pen-Light and Pen-Color experiments consider (12(g)) as source-POMDP and (12(h)) and (12(i)) as the respective target-POMDPs.
VI-E PCA and t-SNE analysis
@@
In Fig. 13 we compare the latent space $\mathcal{Z}$ learned by LAIfO, C-LAIfO, and LAIfO with data augmentation in the Light setting from Table II and Fig. 9. Specifically, we define source and target as the observations generated by an optimal policy in the source and target POMDPs, respectively. Similarly, source random and target random are observations generated by random policies. These observations are processed with the encoder $\phi_{\bm{\delta}}:\mathcal{X}^{d}\to\mathcal{Z}$ trained using the respective algorithms. We perform PCA and t-SNE on this set of latent variables $z_{1:T}$ and plot the first two principal components. The results show that C-LAIfO is the only algorithm capable of filtering out visual distractors and clustering together data points with the same goal-completion information. In Fig. 14, we focus on C-LAIfO and test for generalization to the unseen environment in Fig. 14(c). We train $\phi_{\bm{\delta}}$ on the Full setting from Table II and Fig. 11 and perform PCA and t-SNE as described. In Fig. 14, the Full figures have (14(a))-(14(b)) as source-POMDP and (14(e))-(14(f)) as target-POMDP. Similarly the Full-unseen figures have (14(c))-(14(d)) as source-POMDP and (14(e))-(14(f)) as target-POMDP. The results in the Full-unseen setting match those in the Full experiment, indicating that $\phi_{\bm{\delta}}$ trained on (14(a)) can successfully generalize to (14(c)).
@@
Figure 13: PCA and t-SNE visualizations for the Light experiment in Table II.
@@
Figure 14: PCA and t-SNE for the Full experiment in Table II. Fig. 14(a) and Fig. 14(b) denote the source environment for the Full experiment (left figures in Fig. 14). Fig. 14(c) and Fig. 14(d) denote the source environment for the Full-unseen experiment (right figures in Fig. 14).